solve for the variable in each equation below. show work. provide check for answer. if solution is undefined, state that in answer. thxs
b/2 - b/5 = 6
.2x + .05x = 2
1-.1y > .2
7s-2/3 - 3s-16/6 = 2
8/x-2 = 4

Found 2 solutions by rfer, ewatrrr:
Answer by rfer(16322)

A)
5b/10-2b/10=6
3b/10=6
3b=60
b=20
B)
.25x=2
x=8
C)
-0.1y > 0.2-1
y < -0.8/-0.1
y < 8
you try the others

Answer by ewatrrr(24785)


Hi,
b/2 - b/5 = 6 multiplying each term by 10
5b-2b = 60
b = 20
CHECKING Our Answer***
20/2 - 20/5 = 6
.2x + .05x = 2 multiplying each term by 100
20x + 5x = 200
x = 8
CHECKING Our Answer***
.2*8 + .05*8 = 1.6 + .4 = 2
1-.1y > .2 multiplying each term by 10
10 - y > 2
OR 8 > y
CHECKING Our Answer***
1-.7 >.2
(7s-2)/3 - (3s-16)/6 = 2 multiplying each term by 6
2(7s-2) - (3s-16) = 12
14s - 4 - 3s + 16 = 12
11s = 0
s = 0
CHECKING Our Answer***
-2/3 + 16/6 = -2/3 + 8/3 = 6/3 = 2
8/x-2 = 4/1 Cross multiplying
4(x-2) = 8
x-2 = 2
x = 2
CHECKING Our Answer***
8/(6-4) = 8/2 = 4/1